The phosphorus pentoxide used to produce phosphoric acid for cola soft drinks is prepared by burning phosphorus in oxygen.

P4 + 5O2 ⟶ P4O10

Calculate the percent yield if 9.4 g of P4 reacts with 24.5 g of O2 to yield 4.9 g of P4O10 from the reaction.
Answer #1

First calculate molar masses of reactants and products and their no. of moles.

Molar Mass of P4 = ( 4 x 30.97 ) = 123.88 g / mol

Molar mass of O 2 = ( 2 x 16.00 ) = 32.00 g / mol

Molar mass of P4O10 = ( 4 x 30.97 ) + ( 10 x 16.00 ) = 283.88 g / mol

We know that, No. of moles = Mass / Molar Mass

No. of moles of P4 = 9.4 g / 123.88 g / mol = 0.0759 mol

No. of moles of O 2 = 24.5 g / 32.00 g / mol = 0.766 mol

No. of moles of P4O10 = 4.9 g / 283.88 g / mol = 0.0173 mol

Consider a balanced chemical reaction, P4 + 5 O 2 P4O10

From reaction, stoichiometric ratio of reactants is P4 : O 2 = 1: 5

Provided molar ratio of reactants is P4 : O 2 = 0.0759 : 0.766 = 1: 10.0

Comparing this mole ratio with theoretical mole ratio of reactants it is clear that P4 is limiting reactant. Hence, yield of product will depend on amount of P4.

Calculation of theoretical yield

From reaction , we can write

1 mole P4 1 mole P4O10

123.88 g P4 283.88 g P4O10

9.4 g P4 283.88 x 9.4 / 123.88 g P4O10

21.54 g P4O10

Theoretical yield of product = 21.54 g

Practical yield of product = 4.9 g

% yield of product = [(Practical yield / Theoretical yield) x 100 ]

=( 4.9 g / 21.54 g) x 100

= 22.75 %

ANSWER : % yield of P4O10 = 22.75 %